The probability that San Francisco plays in the next Super Bowl is nine times the probability that they do not play in the next Super Bowl. The probability that San Francisco plays in the next Super Bowl plus the probability that they do not play is 1. What is the probability that San Francisco plays in the next Super Bowl?: The probability that San Francisco plays in the next Super Bowl is nine times the probability that they do not play in the next Super Bowl. The probability that San Francisco plays in the next Super Bowl plus the probability that they do not play is 1. What is the probability that San Francisco plays in the next Super Bowl?
Answer by Fombitz(1756)

P(SF Yes)=9*P(SF No)
P(SF No) = 1/9*P(SF Yes)
P(SF Yes)+P(SF No) = 1
P(SF Yes)+ 1/9*P(SF Yes)=1
P(SF Yes)(1+1/9)=1
P(SF Yes)(10/9)=1
P(SF Yes)=9/10